Find the sum of the first n terms of the following arithmetic sequence.
a1 = 2, d = 2; n = 5.

Answer by rfadrogane(214)

Find the sum of the first n terms of the following arithmetic sequence.
a1 = 2, d = 2; n = 5.
Sol'n:

S = n/2[2a1+(n-1)d]
= (5/2)[2(2)+(5-1)2]
= (5/2)[4+8]
= (5/2)(12)
S = 30 --answer